# Safety through feedback in Constrained RL

### v1

@@ Abstract
In safety-critical RL settings, the inclusion of an additional cost function is often favoured over the arduous task of modifying the reward function to ensure the agent’s safe behaviour. However, designing or evaluating such a cost function can be prohibitively expensive. For instance, in the domain of self-driving, designing a cost function that encompasses all unsafe behaviours (e.g., aggressive lane changes, risky overtakes) is inherently complex, it must also consider all the actors present in the scene making it expensive to evaluate. In such scenarios, the cost function can be learned from feedback collected offline in between training rounds. This feedback can be system generated or elicited from a human observing the training process. Previous approaches have not been able to scale to complex environments and are constrained to receiving feedback at the state level which can be expensive to collect. To this end, we introduce an approach that scales to more complex domains and extends to beyond state-level feedback, thus, reducing the burden on the evaluator. Inferring the cost function in such settings poses challenges, particularly in assigning credit to individual states based on trajectory-level feedback. To address this, we propose a surrogate objective that transforms the problem into a state-level supervised classification task with noisy labels, which can be solved efficiently. Additionally, it is often infeasible to collect feedback on every trajectory generated by the agent, hence, two fundamental questions arise: (1) Which trajectories should be presented to the human? and (2) How many trajectories are necessary for effective learning? To address these questions, we introduce a novelty-based sampling mechanism that selectively involves the evaluator only when the the agent encounters a novel trajectory, and discontinues querying once the trajectories are no longer novel. We showcase the efficiency of our method through experimentation on several benchmark Safety Gymnasium environments and realistic self-driving scenarios. Our method achieves near optimal performance (to when the cost function is known) just from gathering feedback at the trajectory level in many of the domains underscoring the effectiveness and scalability of our approach.
@@ 1 Introduction
Reinforcement Learning (RL) is known to suffer from the problem of reward design, especially in safety-related settings [4]. In such a case, constrained RL settings have emerged as a promising alternative to generate safe policies[3, 14, 32]. This framework introduces an additional cost function to split the task related information (rewards) from the safety related information (costs). In this paper, we address a scenario where the reward function is well-defined, while the cost function remains unknown a priori and requires inference from feedback. This arises in cases where the cost function is expensive to design or evaluate. For instance, consider the development of an autonomous driving system, where the reward function may be defined as the time taken to reach a destination which is easier to define. However, formulating the cost function presents significant challenges. Designing a comprehensive cost function that effectively penalizes all potential unsafe behaviors is non-trivial, and ensuring safety often involves subjective judgments, which can vary based on the individual passenger preferences. Even if one succeeds in devising such a function, it must encapsulate data on all environmental factors, including neighboring vehicles and pedestrians. The evaluation of such a cost function in high-fidelity simulators can be prohibitively expensive.
@@
The feedback can be collected from either a human observer, who monitors the agent’s training process and periodically provides feedback on presented trajectories, or from a system that computes the cost incurred on selected trajectories. Throughout this paper, we use the term evaluator to refer to the entity providing feedback, whether human or system-generated.
@@
Cost inference in Constrained RL settings has gained recent attention. One key thread of research has focused on learning cost from from constraint abiding expert demonstrations [26, 30]. However, these expert demonstrations are not easily available in all settings. For example, consider robotic manipulation tasks where the human and the robot have different morphologies. This paper is focused on the second thread of research in this space, where we generate trajectories and take feedback from an evaluator to infer the cost function. Prior works in this thread make limiting assumptions on the nature of the cost function such as smoothness or linearity [26, 30, 10, 5] and are limited to obtaining feedback at the state level [7] which is expensive to collect from human evaluators. We do not make such assumptions and can take feedback provided at over longer horizons, in some cases the entire trajectory.
@@
Frameworks for learning from feedback must exhibit the following properties to as emphasized in [7, 16, 11]: 1) The feedback must be collected offline in between rounds, since the agent may need to act in real time. 2) The framework should try to minimize the amount of feedback collected. 3) Binary feedback is more suitable as compared to numeric feedback as it is more intuitive to provide for humans. It has also been shown that humans provide less consistent feedback if it is numeric [11]. 4) Each state is assigned a binary cost value, indicating that it is inherently safe or unsafe, which is more intuitive for humans when assessing the safety of policies [29]  [^1] The last two points are specific to human evaluators.
@@
  • Learns safe policies across diverse benchmark safety environments. We also show that the learnt cost function can be used to train an agent with different dynamics/morphology from scratch without collecting additional feedback.
@@ 2 Preliminaries
A Markov Decision Process (MDP) $\mathcal{M}$ is defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma,\mu)$, where $\mathcal{S}$ denotes the set of states, $\mathcal{A}$ is the set of actions, $\mathcal{P}(s^{\prime}|s,a)\in[0,1]$ is the transition probability, $r(s,a)\in\mathbb{R}$ is the reward function, $\gamma\in[0,1]$ is the discount factor and $\mu(s)\in\Delta(\mathcal{S})$ is the initial state distribution. A policy $\pi(.|s)\in\Delta(\mathcal{A})$ is a distribution over the set of valid actions for state $s$. We denote the set of all stationary policies as $\Pi$. A trajectory $\tau=\{(s_{t},a_{t})\}$ denotes the state-action pairs encountered by following $\pi$ in $\mathcal{M}$. We use the short hand $\tau_{i:j}$ to denote a trajectory segment, i.e, the subsequence of $(s_{t},a_{t})$ pairs encountered from timestep $i$ to $j$. The expected value of a function $f$ under $\pi$ as $\mathcal{J}^{f}(\pi)\triangleq E_{\tau\sim\pi}[\sum_{t=0}^{\infty}\gamma^{t}f(s_{t},a_{t})]$. We also employ the shorthand $f(\tau)$ to represent the discounted sum of $f$ along the trajectory $\tau$. The occupancy measure of a policy is define as $\rho(s,a)=E_{\tau\sim\pi}[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{I}[(s_{t},a_{t})=(s,a)]]$, where $\mathbb{I}[.]$ denotes the indicator function. $\rho$ describes the frequency with which a state-action pair is visited by $\pi$.
@@ 4.1 Nature of the Feedback
In the feedback process, first the evaluator is shown the entire trajectory $\tau_{0:T}$. Post this, the trajectory is broken down into contiguous segments $\tau_{i:j}$ of length $k$ and feedback is obtained for each of these segments. Note that the segment length can be set based on the complexity of the environment. For simpler environments, feedback can be collected for the entire trajectory whereas for environments with long horizons and sparse cost violations feedback can be collected for shorter segments. This makes the problem of assigning credit to individual states less challenging. This approach is adopted by other works that rely on human feedback as well [11, 16]. Note that decreasing the segment length comes at a higher cost of obtaining feedback from the evaluator.
The evaluator is tasked with classifying a segment as unsafe if the agent encountered an unsafe state at any point during the segment. The decision was primarily driven by the aim to facilitate consistent feedback from the evaluator. Alternative approaches, such as categorizing the segment as unsafe, if a certain number of unsafe states are visited or leaving the classification upto subjective judgment of the evaluator were deemed prone to generating inconsistent feedback in the case of human evaluators [11].
@@ 4.2 Inferring the Cost Function
Let $P=\{\tau_{i:j},y^{safe}\}$ be the feedback collected from the human, where $y^{safe}=1$ if the segment was labelled safe and vice-versa. We assume that there exists an underlying ground truth cost function $c_{gt}(s,a)\in[0,1]$ based on which the evaluator provides feedback. The probability that a state is safe is defined as $p^{safe}_{gt}(s,a)=\mathbb{I}[c_{gt}(s,a)=0]$. By definition of how the feedback is collected, the probability that a trajectory segment $\tau_{i:j}$ is labelled as safe is given by,
$p^{safe}_{gt}(\tau_{i:j})$ $=\prod_{t=i}^{j}p^{safe}_{gt}(s_{t},a_{t})$ (2)
Now, let $p^{safe}(s,a)$ represent our estimate of $p^{safe}_{gt}(s,a)$ that we intend to learn from the collected feedback. We can optimize $p^{safe}(s,a)$ by minimizing the likelihood loss,
$L^{mle}$ $=E_{(\tau_{i:j},y^{safe})\sim P}\left[{y^{safe}}\log{p^{safe}(\tau_{i:j})}+{(1-y^{safe})}\log{1-p^{safe}(\tau_{i:j})}\right]$
$=E_{(\tau_{i:j},y^{safe})\sim P}\left[{y^{safe}}\log{\prod_{t=i}^{j}p^{safe}(s_{t},a_{t})}+{(1-y^{safe})}\log{(1-\prod_{t=i}^{j}p^{safe}(s_{t},a_{t}))}\right]$
$=E_{(\tau_{i:j},y^{safe})\sim P}\left[{y^{safe}}\sum_{t=i}^{j}\log{p^{safe}(s_{t},a_{t})}+{(1-y^{safe})}\log{(1-\prod_{t=i}^{j}p^{safe}(s_{t},a_{t}))}\right]$ (3)
@@
$L^{sur}$ $=E_{(\tau_{i:j},y^{safe})\sim P}\left[{y^{safe}}\sum_{t=i}^{j}\log{p^{safe}(s_{t},a_{t})}+{(1-y^{safe})}\sum_{t=i}^{j}\log{(1-p^{safe}(s_{t},a_{t}))}\right]$
$=E_{(\tau_{i:j},y^{safe})\sim P}\sum_{t=i}^{j}\sum_{s,a}\mathbb{I}[(s_{t},a_{t})=(s,a)]\Bigg{[}\mathbb{I}[y^{safe}=1]\log{p^{safe}(s,a)}$
@@
$=E_{(s,a)\sim d_{g}}\left[\log{p^{safe}(s,a)}\right]+E_{(s,a)\sim d_{b}}\left[\log{(1-p^{safe}(s,a)})\right]$ (4)
@@
The surrogate loss transforms the objective from the segment level, where issues of collapsing probabilities over lengthy segments arise, to the state level where we do not encounter this problem. Optimizing Eq 4.2 involves breaking down the segments into individual states and assigning each state the label of the segment. Subsequently, states are uniformly sampled at random, and the binary cross-entropy loss is minimized. Thus, $L^{sur}$ represents a binary classification problem where one class contains noisy labels. Every unsafe state is always labelled correctly, as its occurrence forces the evaluator to label the segment unsafe. A safe state is correctly labeled safe when it appears in a segment labeled safe by the evaluator. However, if it occurs in a segment labeled unsafe, it is incorrectly labelled unsafe. We posit that with a sufficient number of samples, the number of correct labels will outweigh the number of incorrect labels for each state.
@@
Proposition 3 highlights that $c^{*}(s,a)$ mislabels some safe states as unsafe if their occurrence in segments labelled unsafe is higher than in segments labelled safe by the human evaluator. We argue that such occurrences are likely to decrease as more data is collected or with a reduced segment length, but recognize that it is guaranteed to be 00 only when the segment length is $1$, i.e, feedback is received at the state-level.
Additionally, note that the bias is greater than or equal to 00, i.e, the expected cost $E_{\pi}[c_{*}(s,a)]$ serves as an upper bound on the true cost incurred by $\pi$. Thus, ensuring that the policy does not exceed the threshold $c_{0}$ on $c_{*}$ guarantees that it adheres to the threshold on $c_{gt}$.
@@ 4.3 Efficient Subsampling of Trajectories
To reduce the burden on the evaluator (and reduce the cost of feedback), we present a subset of the trajectories collected by the policy for feedback. The common approach is to break the problem into two parts: (1) define a schedule $N_{queries}(i)$ that determines the number of trajectories to be shown to the user at the end of each data collection round $i$. Subsequently, $N_{queries}(i)$ trajectories are sampled from those collected by the policy at data collection round $i$. While ideally, we aim to select the subset of trajectories maximizing the expected value of information [17], this task is computationally intractable [2]. To address this challenge, various sampling methods have been employed, seeking to maximize a surrogate measure of this value. Among these, uncertainty sampling stands out as the most prominent approach, wherein trajectories are sampled based on the estimator’s uncertainty about their predictions [11, 16, 18]. However, quantifying this uncertainty is non-trivial given the lack of calibration in neural network predictions. To address this challenge, ensemble methods are frequently employed where the disagreement among the models is used as an uncertainty measure. However, the training of $n$ distinct neural networks can exact substantial resource costs, prompting consideration for alternative approaches
In light of this, we introduce a new form of uncertainty sampling called novelty sampling. With novelty sampling, we gather all the novel trajectories after each round and present them to the evaluator for feedback. Formally, we define a state as novel if its density in the feedback data collected so far $d(s)=\sum_{a}d(s,a)$ is 00. A trajectory is deemed novel if it comprises of at least $e$ novel states. This can be interpreted as as ensuring the edit distance (a popular trajectory distance measure [25]) between the trajectory and previously seen trajectories is greater than $e$. We do consider novelty for state-action pairs as we found that extending to this case adversely impacted the performance.
The central notion is that the model is prone to errors on novel states- those it has not encountered during training. This arises because the policy evolves over time, venturing into states that were not previously encountered during data collection rounds. Hence, this sampling strategy effectively reduces the epistemic uncertainty of the model- error stemming from a lack of training data. Furthermore, this sampling method offers the advantage of implicitly establishing a decreasing querying schedule as novelty of trajectories reduces over time as shown in Figure 6 in the Appendix.

### v2

@@ Abstract
In safety-critical RL settings, the inclusion of an additional cost function is often favoured over the arduous task of modifying the reward function to ensure the agent’s safe behaviour. However, designing or evaluating such a cost function can be prohibitively expensive. For instance, in the domain of self-driving, designing a cost function that encompasses all unsafe behaviours (e.g., aggressive lane changes, risky overtakes) is inherently complex, it must also consider all the actors present in the scene making it expensive to evaluate. In such scenarios, the cost function can be learned from feedback collected offline in between training rounds. This feedback can be system generated or elicited from a human observing the training process. Previous approaches have not been able to scale to complex environments and are constrained to receiving feedback at the state level which can be expensive to collect. To this end, we introduce an approach that scales to more complex domains and extends beyond state-level feedback, thus, reducing the burden on the evaluator. Inferring the cost function in such settings poses challenges, particularly in assigning credit to individual states based on trajectory-level feedback. To address this, we propose a surrogate objective that transforms the problem into a state-level supervised classification task with noisy labels, which can be solved efficiently. Additionally, it is often infeasible to collect feedback for every trajectory generated by the agent, hence, two fundamental questions arise: (1) Which trajectories should be presented to the human? and (2) How many trajectories are necessary for effective learning? To address these questions, we introduce a novelty-based sampling mechanism that selectively involves the evaluator only when the the agent encounters a novel trajectory, and discontinues querying once the trajectories are no longer novel. We showcase the efficiency of our method through experimentation on several benchmark Safety Gymnasium environments and realistic self-driving scenarios. Our method demonstrates near-optimal performance, comparable to when the cost function is known, by relying solely on trajectory-level feedback across multiple domains. This highlights both the effectiveness and scalability of our approach. The code to replicate these results can be found at https://github.com/shshnkreddy/RLSF
@@ 1 Introduction
Reinforcement Learning (RL) is known to suffer from the problem of reward design, especially in safety-related settings [4]. In such a case, constrained RL settings have emerged as a promising alternative to generate safe policies[3, 16, 34]. This framework introduces an additional cost function to split the task related information (rewards) from the safety related information (costs). In this paper, we address a scenario where the reward function is well-defined, while the cost function remains unknown a priori and requires inference from feedback. This arises in cases where the cost function is expensive to design or evaluate. For instance, consider the development of an autonomous driving system, where the reward function may be defined as the time taken to reach a destination which is easy to define. However, formulating the cost function presents significant challenges. Designing a comprehensive cost function that effectively penalizes all potential unsafe behaviors is non-trivial, and ensuring safety often involves subjective judgments, which can vary based on individual preferences. Even if one succeeds in devising such a function, it must account for all environmental factors, such as neighboring vehicles and pedestrians. The evaluation of such a cost function in high-fidelity simulators can be prohibitively expensive.
@@
Feedback can be collected from either a human observer, who monitors the agent’s training process and periodically provides feedback on presented trajectories, or from a system that computes the cost incurred on selected trajectories. Throughout this paper, we use the term evaluator to refer to the entity providing feedback, whether human or system-generated.
@@
Cost inference in Constrained RL settings has gained recent attention. One key thread of research has focused on learning cost from from constraint abiding expert demonstrations [28, 32]. However, these expert demonstrations are not easily available in all settings. For example, consider robotic manipulation tasks where the human and the robot have different morphologies. This paper focuses on the second line of research in this area, where we generate trajectories and gather feedback from an evaluator to infer the underlying cost function. Prior works in this thread make limiting assumptions on the nature of the cost function such as smoothness or linearity [28, 32, 11, 5] and are limited to obtaining feedback at the state level [8] which is expensive to collect from human evaluators. We do not make such assumptions and can take feedback provided at over longer horizons, in some cases the entire trajectory.
@@
Frameworks for learning from feedback must exhibit the following properties as emphasized in [8, 18, 12]: 1) Feedback must be collected offline in between rounds, since the agent may need to act in real time. 2) The amount of feedback collected must be minimized. 3) Binary feedback is more suitable as compared to numeric feedback as it is more intuitive to provide for humans. It has also been shown that humans provide less consistent feedback if it is numeric [12]. 4) Each state is assigned a binary cost value, indicating that it is inherently safe or unsafe, which is more intuitive for humans when assessing the safety of policies [31]  [^1] The last two points are specific to human evaluators.
@@
  • Learns safe policies across diverse benchmark safety environments. We also show that the learnt cost function can be transferred to train an agent with different dynamics/morphology from scratch without collecting additional feedback.
@@ 2 Preliminaries
A Markov Decision Process (MDP) $\mathcal{M}$ is defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma,\mu)$, where $\mathcal{S}$ denotes the set of states, $\mathcal{A}$ is the set of actions, $\mathcal{P}(s^{\prime}|s,a)\in[0,1]$ is the transition probability, $r(s,a)\in\mathbb{R}$ is the reward function, $\gamma\in[0,1]$ is the discount factor and $\mu(s)\in\Delta(\mathcal{S})$ is the initial state distribution. A policy $\pi(.|s)\in\Delta(\mathcal{A})$ is a distribution over the set of valid actions for state $s$. We denote the set of all stationary policies as $\Pi$. A trajectory $\tau=\{(s_{t},a_{t})\}$ denotes the state-action pairs encountered by executing $\pi$ in $\mathcal{M}$. We use the short hand $\tau_{i:j}$ to denote a trajectory segment, i.e, the subsequence of $(s_{t},a_{t})$ pairs encountered from timestep $i$ to $j$. The expected value of a function $f$ under $\pi$ as $\mathcal{J}^{f}(\pi)\triangleq E_{\tau\sim\pi}[\sum_{t=0}^{\infty}\gamma^{t}f(s_{t},a_{t})]$. We also employ the shorthand $f(\tau)$ to represent the discounted sum of $f$ along the trajectory $\tau$. The occupancy measure of a policy is define as $\rho(s,a)=E_{\tau\sim\pi}[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{I}[(s_{t},a_{t})=(s,a)]]$, where $\mathbb{I}[.]$ denotes the indicator function. $\rho$ describes the frequency with which a state-action pair is visited by $\pi$.
@@ 4.1 Nature of the Feedback
In the feedback process, the evaluator is first presented with the entire trajectory $\tau_{0:T}$. Afterward, the trajectory is divided into contiguous segments $\tau_{i:j}$ of length $k$, and feedback is collected for each segment. The segment length can be adjusted based on the complexity of the environment: in simpler environments, feedback can be gathered for the entire trajectory, whereas for environments with long horizons and sparse cost violations, shorter segments may be used. This approach simplifies the challenge of assigning credit to individual states. Similar methods have been adopted in other works that rely on human feedback [12, 18]. However, reducing segment length comes at an increased cost of obtaining feedback from the evaluator.
The evaluator is tasked with classifying a segment as unsafe if the agent encounters an unsafe state at any point within the segment. This decision was made to ensure consistent feedback from the evaluator. Alternative approaches—such as marking a segment unsafe based on the number of unsafe states visited or leaving the classification to the evaluator’s subjective judgment are more prone to generating inconsistent feedback in the case of human evaluators [12].
@@ 4.2 Inferring the Cost Function
Let $P=\{\tau_{i:j},y^{safe}\}$ represent the feedback collected from the evaluator, where $y^{safe}=1$ if the segment was labelled safe and $y^{safe}=0$ otherwise. We assume there exists an underlying ground truth cost function $c_{gt}(s,a)\in[0,1]$ based on which the evaluator provides feedback. Then, the probability that a state is safe is defined as $p^{safe}_{gt}(s,a)=\mathbb{I}[c_{gt}(s,a)=0]$.
Now, let $p^{safe}(s,a)$ represent our estimate of $p^{safe}_{gt}(s,a)$ that we intend to estimate from the collected feedback. Then, by definition of how the feedback is collected, the probability that a trajectory segment $\tau_{i:j}$ is labelled as safe is given by,
$p^{safe}(\tau_{i:j})$ $=\prod_{t=i}^{j}p^{safe}(s_{t},a_{t})$ (2)
We can infer $p^{safe}(s,a)$ by minimizing the likelihood loss,
$L^{mle}$ $=-E_{(\tau_{i:j},y^{safe})\sim P}\left[{y^{safe}}\log{p^{safe}(\tau_{i:j})}+{(1-y^{safe})}\log{1-p^{safe}(\tau_{i:j})}\right]$
$=-E_{(\tau_{i:j},y^{safe})\sim P}\left[{y^{safe}}\sum_{t=i}^{j}\log{p^{safe}(s_{t},a_{t})}+{(1-y^{safe})}\log{(1-\prod_{t=i}^{j}p^{safe}(s_{t},a_{t}))}\right]$ (3)
@@
$L^{sur}$ $=-E_{(\tau_{i:j},y^{safe})\sim P}\left[{y^{safe}}\sum_{t=i}^{j}\log{p^{safe}(s_{t},a_{t})}+{(1-y^{safe})}\sum_{t=i}^{j}\log{(1-p^{safe}(s_{t},a_{t}))}\right]$
$=-E_{(\tau_{i:j},y^{safe})\sim P}\sum_{t=i}^{j}\sum_{s,a}\mathbb{I}[(s_{t},a_{t})=(s,a)]\Bigg{[}\mathbb{I}[y^{safe}=1]\log{p^{safe}(s,a)}$
@@
$=-\left[E_{(s,a)\sim d_{g}}\log{p^{safe}(s,a)}+E_{(s,a)\sim d_{b}}\log{(1-p^{safe}(s,a)})\right]$ (4)
@@
The surrogate loss reformulates the objective from the segment level—where collapsing probabilities over long segments can be problematic—to the state level, where this issue does not occur. Optimizing Eq 4.2 involves breaking down the segments into individual states and assigning each state the label of the segment. Subsequently, states are uniformly sampled at random, and the binary cross-entropy loss is minimized. Consequently, $L^{sur}$ represents a binary classification problem in which one class contains noisy labels. While unsafe states are always accurately labeled—since their presence necessitates an unsafe classification for the segment—a safe state may receive conflicting labels based on the status of the segment it belongs to. Thus, with a sufficient number of samples, we believe it becomes feasible to reliably differentiate between safe and unsafe states.
@@
Proposition 3 illustrates that $c^{*}(s,a)$ misclassifies certain safe states as unsafe when their frequency in segments labeled unsafe exceeds that in segments labeled safe by the evaluator. We contend that this misclassification is likely to diminish with increased data collection or shorter segment lengths. However, it is important to note that this misclassification is guaranteed to be zero only when the segment length is reduced to one, meaning feedback is provided at the state level.
Additionally, note that the bias is non-negative, meaning that the expected cost $E_{\pi}[c_{*}(s,a)]$ acts as an upper bound on the true cost incurred by $\pi$. Therefore, ensuring that the policy does not exceed the threshold $c_{max}$ on $c_{*}$ guarantees that it adheres to the threshold on $c_{gt}$.
@@ 4.3 Efficient Subsampling of Trajectories
To reduce the burden on the evaluator and minimize the cost of feedback, we present a subset of the trajectories collected by the policy for feedback. The common approach is to break the problem into two parts: (1) define a schedule $N_{queries}(i)$ that determines the number of trajectories to be shown to the user at the end of each data collection round $i$. Subsequently, $N_{queries}(i)$ trajectories are sampled from those collected by the policy at data collection round $i$. While the ideal goal is to sample a subset of trajectories that maximizes the expected value of information [19], achieving this is computationally intractable [2]. To address this challenge, various sampling methods have been employed, seeking to maximize a surrogate measure of this value. Among these, uncertainty sampling stands out as the most prominent approach, wherein trajectories are sampled based on the estimator’s uncertainty about their predictions [12, 18, 20]. However, quantifying the uncertainty is challenging given the lack of calibration in neural network predictions. To address this challenge, ensemble methods are frequently employed where the disagreement among the models is used as an uncertainty measure. However, the training of $n$ distinct neural networks can exact substantial resource costs, prompting consideration for alternative approaches.
In light of this, we introduce a new form of uncertainty sampling called novelty sampling. With novelty sampling, we gather all the novel trajectories after each round and present them to the evaluator for feedback. Formally, we define a state as novel if its density in the feedback data collected so far $d(s)=\sum_{a}d(s,a)$ is 00. A trajectory is deemed novel if it comprises of at least $e$ novel states. This can be interpreted as ensuring that the edit distance—a well-known measure of trajectory distance [27]—between the current trajectory and previously seen trajectories exceeds a threshold $e$. We do not consider novelty for state-action pairs as we found that extending to this case adversely impacted the performance.
The central notion is that the model is prone to errors on novel states- those it has not encountered during training. This arises because the policy evolves over time, venturing into states that were not previously encountered during data collection rounds. Therefore, this sampling strategy effectively reduces the epistemic uncertainty of the model—error arising from insufficient training data—thereby making it a form of uncertainty sampling. Furthermore, this sampling method offers the advantage of implicitly establishing a decreasing querying schedule as novelty of trajectories reduces over time as shown in Figure 6 in the Appendix.